OpenStudy (anonymous):

find a second degree polynomial with real coefficients that has a root at x=2i

OpenStudy (anonymous):

If \(2i\) is a root then so is \(-2i\).

OpenStudy (zehanz):

Remember that if such a polynomial has root 2i then it also has -2i as a root. Now you can write a factored form of it as follows: (x-2i)(x+2i). Expand (FOIL) and there is your polynomial. In fact, any number a (nonzero) can be put in front of the brackets: a(x-2i)(x+2i). There are infinitely many functions. Leaving a=1 gives you the simplest one...

OpenStudy (anonymous):

I cant foil an x w 2i

OpenStudy (zehanz):

Yes you can! it is of the form (a+b)(a-b)=a²-b² (remember that one?) So (x-2i)(x+2i) = x² - (4i²) = x²-4*-1=x²+4

OpenStudy (anonymous):

so thatd b my final answer?

OpenStudy (zehanz):

Yes, that is the polynomial of degree 2 with 2i as a root (and also -2i).

OpenStudy (anonymous):

so my final answer is x^2+4

OpenStudy (zehanz):

Yes! Final, definite, 100%!